# Discovery of Critical Thresholds in Mixed Exposures and Estimation of Policy Intervention Effects using Targeted Learning

### v2

@@ 4.1 G-Computation Decision Tree
We introduce a novel decision tree algorithm developed to approximate the oracle parameter. This parameter, denoted as $\underset{\mathcal{A}}{\mathrm{argmin}}\ E[E[Y|\mathcal{A}=1,W]]$, represents the region $\mathcal{A}$ that minimizes the expected outcome for a given set of covariates $W$. Our approach is tailored to meet several key criteria: computational efficiency, interpretability, and implicit regularization to avoid overly granular and statistically insignificant regions. The core of our method is a regression tree that optimizes a simple plug-in estimator for the policy effect ($E[Y|\tilde{A}=1,W]$) of a given region, based on a Random Forest or a similar straightforward model. The decision tree is constructed to search for the minimum/maximum region using this plug-in estimator, greedily evaluating splits in the exposure variables $A$ and selecting the region that maximizes/minimizes the estimated average outcome in the newly formed regions.
@@
In the partitioning criterion phase, the algorithm explores all binary splits across the range of each exposure variable to identify the optimal partitioning strategy. Consider a scenario where we have five exposure variables, A1 through A5. The algorithm examines the unique values for each exposure variable. For example, for A1, it evaluates the potential splits at A1 < 1, A1 < 2, and so forth, up to A1 < m, where m represents the maximum observed value of A1. At each of these split points, the algorithm constructs a binary indicator defining the region of interest and employs the plug-in estimator. This estimator is trained to assess the mean outcome under the hypothetical situation where all individuals are ‘forced’ into that region. The algorithm’s objective at each split point is to minimize/maximize the average predicted outcome within the newly formed child nodes, relative to the parent node. This process of finding the optimal split is applied to each of the exposure variables. For example, if the algorithm identifies that splitting at A5 < 5 yields the minimum average outcome, this split becomes the root node of our decision tree. Subsequently, the algorithm partitions the data into two segments based on the chosen split, resulting in a left node (A5 $<$ 5) and a right node (A5 $\geq$ 5). The recursive process is then applied to each of these data regions demarcated by the node, iterating through all exposure variables again to find further splits that minimize the average outcome. For example, within the left node, the algorithm might test combinations such as A5 $<$ 5 &\&& A1 $<$ 2, etc., each time training the plug-in estimator and evaluating the mean outcome for the defined region. This recursive partitioning continues until two stopping criteria are met: it either reaches the pre-specified maximum depth of the tree, or it arrives at a point where no further split can yield an average outcome lower/greater than that of the parent node.
@@ 5.1 Pooled TMLE
Our substitution estimator, denoted by $\bar{Q}_{n}$, approximates the true conditional mean $\bar{Q}_{0}$ by plugging in the empirical distribution for each observation. The estimator is operationalized via a Super Learner algorithm followed by TMLE, and its cross-validated variant is expressed as $\Psi(Q_{P_{n_{k}}}^{\star})=\frac{1}{V}\sum_{v=1}^{V}\{\bar{Q}_{n_{-k}}^{\star}(\mathcal{A}_{n_{-k}}=1,W_{v})-Y_{v}\}$. Where $Q_{n_{-k}}^{\star}$ is the TMLE updated expectation of the outcome if everyone were exposed to levels of exposure within the data-adaptively determined region. The cross-estimation approach not only utilizes the full dataset for variance estimation, resulting in tighter confidence intervals, but also enables the derivation of fold-specific and pooled estimates.
@@
That is, this pooled parameter represents the ARE for the determined maxmizing/minimizing region in the exposure space. Of course, the actual regions determined in folds can vary, and thus this acts as an omnibus test, which can be used to evaluate fold-specific results. A significant pooled TMLE ARE for the oracle region can be interpreted as ’There is a region in the exposure space where the expected outcome is significantly different compared to the average outcome.’ The analyst can then evaluate the fold-specific results to determine what regions make up this oracle estimate. Of course, interpretability of the pooled oracle estimate is reliant on regions in the same sets of exposures that are found across the folds. Otherwise, this pooled estimate allows the analyst to understand if there is generally "signal" in the data for a region or set of regions that collectively differ in expected outcomes compared to the average. This dual presentation allows for the evaluation of variability across folds, offering insight into the stability of the pooled ARE. In particular, when exposure regions $\mathcal{A}$ exhibit high variability between folds, fold-specific results provide an essential interpretive counterbalance to the pooled estimates.
We also provide TMLE updated estimates for the region for the same sets of exposures. For example, if a region was found using thresholds for exposures A1 and A3 in 7 out of 10 folds, we pool the results for these folds for this exposure set. Of course, because 3 folds did not include these exposures for the oracle region, we need to incorporate an inferred null estimate and weigh our pooled estimate incorporating these null results. We discuss this weighted pooling in the next section.
5.2 Pooled TMLE Incorporating Null Results
It is possible that the region with the minimum or maximum expected outcome is not found in every fold. This makes the pooled result difficult to interpret, as these estimates do not incorporate information from folds with null results. As such, in refining our approach to pooled TMLE estimation, our objective is to integrate both the non-null estimates from the estimation samples and the null values. We propose a weighted average approach that combines our original pooled TMLE estimate with the implied null value, which we define as 0. Specifically, let $n_{0}$ denote the number of folds where the null is implied, and let $n_{1}$ represent the number of folds with non-null estimates. The variance of the "null" estimate, $var(\Psi_{\text{null}})$, is estimated using the variance of the original pooled estimates, $var(\hat{\Psi}_{\text{pooled}})$, as $var(\Psi_{\text{null}})=\left(var(\hat{\Psi}_{\text{pooled}})\right)\cdot\frac{n_{1}}{n_{0}}$. We then calculate the weighted average of the original pooled estimate and the null value: $\hat{\Psi}_{\text{pooled}}^{*}=\frac{w_{1}\cdot\hat{\Psi}_{\text{pooled}}+w_{0}\cdot\Psi_{\text{null}}}{w_{0}+w_{1}}$, where $w_{1}=\frac{1}{var(\hat{\Psi}_{\text{pooled}})}$ and $w_{0}=\frac{1}{var(\Psi_{\text{null}})}$. The variance of this new pooled estimate is given by $var(\hat{\Psi}_{\text{pooled}}^{*})=\frac{1}{w_{1}+w_{0}}$. This approach ensures that our pooled TMLE estimate reflects both the explicit estimates from folds where the maximizing/minimizing region was selected and the implicit null values from the remaining folds. In the case where the oracle region in all folds is contained in the same exposure set, we can average the rules to create an interpretable rule for the pooled TMLE estimate. For example, given three partitions like $X_{1}<2$ $X_{2}>5$, $X_{1}<2.3$ $X_{2}>5.2$, and $X_{1}<1.9$ $X_{2}>5.3$, the average rule becomes $X_{1}<2.06(1.9,2.3)$ $X_{2}>5.16(5,5.3)$. The numbers in parentheses are the lower and upper bounds of the threshold for the oracle region, which give an estimate of uncertainty for the thresholds. Overall, interpretability of the estimate can be straightforward if there is consistency in where the oracle region is in the exposure space across folds. If the same exposures contain the oracle region across all the folds, the analyst can simply interpret the pooled TMLE ARE as the estimate for the average oracle region. However, if there is variability, the pooled TMLE acts as a sanity check for any set of regions, leading to a significant difference in the average outcome. If so, the analyst can then look at the k-fold specific results, which also have valid inference, and the pooled estimates within each exposure set while also reporting the number of folds a specific exposure set contained the oracle region.
@@ 6 Simulations
In this section, we demonstrate using simulations that our approach identifies the correct exposure region that maximizes/minimizes conditional mean and estimates the correct PIE target parameter built into a DGP for this region.
@@ 6.1.1 Two-Dimensional Exposure Simulations
This DGP has the following characteristics, $O=(W,A,Y)$. $W$ are three baseline covariates, $W_{1}\sim\mathcal{N}(\mu=37,\sigma=3),W_{2}\sim\mathcal{N}(\mu=20,\sigma=1),W_{3}\sim\mathcal{B}(\mu=0.5)$. Here $\mathcal{B}$ is a Bernoulli distribution, and $\mathcal{N}$ is normal. These distributions and values were chosen to represent a study with covariates for age, BMI, and sex. Our generated exposures were also created to represent two chemical exposures quantized into five discrete levels. The values and range of the outcomes were chosen to represent common environmental health outcomes, such as telomere length or epigenetic expression. We are interested in sampling observations into a 2-dimensional exposure grid. We use discrete exposures to facilitate calculation of the ground truth effects without numerical integration. Here, a $5\times 5$ grid is based on combinations of two discrete exposure levels with values 1-5. We want the number of observations in each of these cells to be affected by covariates. To do this, we define a conditional categorical distribution $P\{(A_{1},A_{2})=(a_{1},a_{2})|W=w\}$ and sample from it. $P\{(A_{1},A_{2})=(a_{k},a_{l})|W\}=\frac{e^{W^{\top}\beta_{k,l}}}{1+\sum_{k,l}e^{W^{\top}\beta_{k,l}}}$. Here, the $\beta$’s attached to each covariate were drawn from a normal distribution with means 0.3, 0.4, 0.5 and 0.5, respectively, all with a standard deviation of 2. This then gives us 25 unique exposure regions with densities dependent on the covariates. We then want to assign an outcome to each of these regions based on the main effects and interactions between exposures. We generate the outcome as: $Y=0.2A_{1}^{2}+0.5A_{1}A_{2}+0.5A_{2}^{2}+0.2*W_{1}+0.4*W_{3}+\mathcal{N}(0,0.1)$. This indicates that there is a slightly weaker squared effect for $A_{2}$ relative to $A_{1}$ and a strong interaction between the exposures and confounding due to $W_{1}$ and $W_{2}$.
@@
Therefore, to approximate the ARE of a region $\mathcal{A}$ to arbitrary precision we compute $\text{ARE}(\mathcal{A})=\frac{1}{b}\sum_{i}^{b}Q_{\mathcal{A}}(1,W_{i})-Y_{i}$ using a large $b$. We repeat this for each of the $(6^{2})^{2}$ possible rectangular regions to determine which has the optimal ARE.
@@ 6.2 Evaluating Performance
For each iteration, we calculate metrics for bias, variance, MSE, CI coverage, and confusion table metrics for the true minimal region compared to the estimated region. We also calculate the bias for the estimated mean in the region found in our recursive partitioning algorithm to find the minimum region. For each type of estimate (pooled TMLE, k-fold specific TMLE estimates, and inverse variance, if applicable), we have bias when comparing our estimate to 1. The ARE based on the true region in the DGP that minimizes the outcome mean 2. The ARE when the data-adaptively determined region is applied to the population. Therefore, when comparing to the true "oracle" region ARE we have $\psi^{0}_{\text{pooled tmle MSE}}$: This is the MSE of the pooled TMLE ARE compared to the ground-truth ARE for the true oracle region built into the DGP which minimizes the mean adjusted outcome and $\psi^{0}_{\text{mean v-fold tmle MSE}}$: This is the MSE of the mean k-fold specific AREs compared to the ground-truth ARE for the true region built into the DGP.
@@ 6.4.1 CVtreeMLE Algorithm Identifies the Oracle Region with Minimum Expected Outcome
First, we describe the results to identify the true region built into the DGP. It is obviously necessary that this converges to the truth as the sample size increases in order for the estimates $\psi^{0}$ to be asymptotically unbiased. Overall, we find that our tree algorithm identifies the true region in the DGP and therefore provides results that have a minimum outcome for treatment policies. Our recursive algorithm shows very good identification of the true outcome minimizing region even in low sample sizes. At sample size 300 the average false positives were 0.053, at 500, 0.0005, and at 750 onward 0. The average true positives were 100% across all sample sizes. Our recursive partitioning algorithm correctly identifies the interaction between both exposures as the minimizing region (as opposed to marginal impacts) in 100% of folds in all sample sizes. As such, we discuss the pooled TMLE results which are the same as the inverse variance results, since in no fold were there null results. Figure 1 shows the bias in comparing the estimated average in the estimated minimum region of the decision tree to that of the true oracle average in the oracle minimizing region. From this figure it can be seen that we have asymptotic convergence to the true minimal average at a rate faster than $\sqrt{n}$ for this simulation.

### v3

@@ 4.1 Targeted Decision Trees
We introduce a novel decision tree algorithm developed to approximate the oracle parameter. This parameter, denoted as $\underset{\mathcal{A}}{\mathrm{argmin}}\ E[E[Y|T=1,W]]$, represents the region $\mathcal{A}$ that minimizes the expected outcome for a given set of covariates $W$. Our approach is tailored to meet several key criteria: computational efficiency, interpretability, and implicit regularization to avoid overly granular and statistically insignificant regions. The core of our method is a regression tree that optimizes a simple plug-in estimator for the policy effect ($E[Y|T=1,W]$) of a given axis-aligned rectangular region, based on a Random Forest or a similar straightforward model paired with targeted learning. The decision tree is constructed to search for the minimum/maximum region using this plug-in estimator, greedily evaluating splits in the exposure variables $A$ and selecting the region that maximizes/minimizes the estimated average outcome in the newly formed regions.
Simply using g-computation to estimate the conditional mean outcome given a region and covariates may run into issues due to positivity. In areas of the exposure space that are very unlikely for certain $w$, our estimate of the mean outcome may behave in strange ways, especially when using nonparametric estimators. Model extrapolation could lead to very wild estimates in certain regions, but the data adaptive region might be selected specifically because they have large (or small) estimates of the mean outcome. To embed in our algorithm this restriction of only selecting regions with support, we incorporate two restrictions 1. Partitioning occurs only if the conditional probability of being in the region is bounded away from 0 and 1 and 2. we debias our initial plug-in estimates of the conditional mean in the left and right regions of the tree using targeted learning and determine if there is a significant difference compared to the parent node mean by calculating the influence function for the population intervention effect, comparing the region means to the parent mean and getting estimates of variance for p-value calculations using the influence function. We give a detailed pseudocode of the algorithm in Algorithm 1, a brief synopsis is given below.
@@
In the partitioning criterion phase, the algorithm explores all binary splits across the range of each exposure variable to identify the optimal partitioning strategy. Consider a scenario where we have five exposure variables, A1 through A5. The algorithm examines the unique values for each exposure variable. For example, for A1, it evaluates the potential splits at A1 < 1, A1 < 2, and so forth, up to A1 < m, where m represents the maximum observed value of A1. At each of these split points, the algorithm constructs a binary indicator defining the region of interest and employs the plug-in estimator. This estimator is trained to assess the mean outcome under the hypothetical situation where all individuals are ‘forced’ into that region. The regions are to the left and right of the split point. For each, we debias the initial estimate using targeted learning and assess the significance of the region mean relative to the parent node mean by estimating the influence function for the population intervention effect, comparing the region node means to the parent node mean.
The algorithm’s objective at each split point is to minimize/maximize the average predicted outcome within the newly formed child nodes, relative to the parent node. This process of finding the optimal split is applied to each of the exposure variables. For example, if the algorithm identifies that splitting at A5 < 5 yields the minimum average outcome, this split becomes the root node of our decision tree. Subsequently, the algorithm partitions the data into two segments based on the chosen split, resulting in a left node (A5 $<$ 5) and a right node (A5 $\geq$ 5). The recursive process is then applied to each of these data regions demarcated by the node, iterating through all exposure variables again to find further splits that minimize the average outcome. For example, within the left node, the algorithm might test combinations such as A5 $<$ 5 &\&& A1 $<$ 2, etc., each time training the plug-in estimator and evaluating the mean outcome for the defined region. This recursive partitioning continues until two stopping criteria are met: it either reaches the pre-specified maximum depth of the tree, or it arrives at a point where no further split can yield an average outcome significantly lower/greater than that of the parent node.
Algorithm 1 Recursive Partitioning Decision Tree using TMLE
1:Input: Data $D$, exposures $\mathbf{A}$, covariates $W$, outcome $Y$, max depth $d_{\text{max}}$, min or max
2:Output: Rule that demarcates the region optimizing $E[Y|\mathcal{A}^{*}=1,\boldsymbol{A},W]$
3:procedure TargetPart($D$, $\mathbf{A}$, $W$, $Y$, depth=0)
4:     if depth $=d_{\text{max}}$ then
5:         return $best\_split$triangleright Stop recursion
6:     end if
7:     Init $best\_split\leftarrow\text{Null}$, $p_{\text{best}}\leftarrow 1$, $min\_avg\leftarrow\infty$, $max\_avg\leftarrow-\infty$
8:     for $a\in\mathbf{A}$ do
9:         for $s$ in unique values of $a$ do
10:              $\mathcal{A}^{*}\leftarrow(D[a]\leq s)$
11:              Fit $P(\mathcal{A}^{*}|W)$; get propensities $p$
12:              if any p < 0.001 or $p$ > 0.99 then $\triangleright$ continue
13:              end if
14:              Calculate $H(a,w)$ from $p$
15:              Fit outcome model $E[Y|\mathcal{A}^{*},\boldsymbol{A},W]$; predict for $\mathcal{A}^{*}=1,0$
16:              Update predictions (TMLE), compare left and right region averages to parent average, $\theta_{\text{left}},\theta_{\text{right}}$
17:              Calculate SEs from IC, p-values; update $best\_split$ if p-value $<$ $p_{\text{best}}$
18:         end for
19:     end for
20:     if $best\_split$ then
21:         $D_{\text{left}}\leftarrow D[\mathcal{A}^{*}=1]$, $D_{\text{right}}\leftarrow D[\mathcal{A}^{*}=0]$
22:         Recursively apply TargetPart(data subset) based on $best\_split$
23:     end if
24:end procedure
25:Initiate TargetPart($D$, $\mathbf{A}$, $W$, $Y$, 0)
@@ 5.1 Pooled TMLE
Our substitution estimator, denoted by $\bar{Q}_{n}$, approximates the true conditional mean $\bar{Q}_{0}$ by plugging in the empirical distribution for each observation. The estimator is operationalized via a Super Learner algorithm followed by TMLE, and its cross-validated variant is expressed as
$\Psi(Q_{P_{n_{k}}}^{\star})=\frac{1}{V}\sum_{v=1}^{V}\{\bar{Q}_{n_{-k}}^{\star}(T_{n_{-k}}=1,W_{v})-Y_{v}\}$
@@
Where $Q_{n_{-k}}^{\star}$ is the TMLE updated expectation of the outcome if everyone were exposed to levels of exposure within the data-adaptively determined region. The cross-estimation approach not only utilizes the full dataset for variance estimation, resulting in tighter confidence intervals, but also enables the derivation of fold-specific and pooled estimates.
@@
That is, this pooled parameter represents the ARE for the determined maxmizing/minimizing region in the exposure space. Of course, the actual regions determined in folds can vary, and thus this acts as an omnibus test, which can be used to evaluate fold-specific results. A significant pooled TMLE ARE for the oracle region can be interpreted as “there is a region in the exposure space where the expected outcome is significantly different compared to the average outcome”. The analyst can then evaluate the fold-specific results to determine what regions make up this oracle estimate. Of course, interpretability of the pooled oracle estimate is reliant on regions in the same sets of exposures that are found across the folds. Otherwise, this pooled estimate allows the analyst to understand if there is generally "signal" in the data for a region or set of regions that collectively differ in expected outcomes compared to the average. This dual presentation allows for the evaluation of variability across folds, offering insight into the stability of the pooled ARE. In particular, when exposure regions $\mathcal{A}$ exhibit high variability between folds, fold-specific results provide an essential interpretive counterbalance to the pooled estimates.
5.2 Theoretical Properties of Region Holding and ARE Convergence
Theoretical results for data-adaptive parameters are shown in Hubbard et al. [2016] which established the asymptotic properties of these parameters under a framework that leverages cross-validation through data partitioning into $V$ equally sized sub-samples.
The estimator developed for each partition is not only consistent but also asymptotically normal due to the independence of the partitioned samples and the efficient estimator applied within each partition. They demonstrate that the average of these estimators, denoted as $\hat{\psi}_{n}$, converges in distribution to a normal distribution centered at the true data-adaptive parameter $\psi_{0}$, with a variance that can be consistently estimated. This convergence is underpinned by the central limit theorem, which applies due to the independent and identically distributed nature of the estimators across partitions.
Furthermore, it is crucial to address the role of the restriction on the size of the exposure region in our case, based on the positivity assumption. This assumption ensures that each exposure level within the region has a non-negligible probability of occurrence, which is critical for maintaining the validity of statistical inferences as the sample size increases. Without this assumption, as sample sizes grow, the selected region might progressively shrink, potentially invalidating the model due to the violation of the positivity condition. Hence, maintaining a sufficiently large region is vital for ensuring that the estimators converge appropriately and that the asymptotic properties hold, thereby avoiding biases that could arise from increasingly narrow exposure definitions.
@@ 6 Simulations
In this section, we demonstrate using simulations that our approach identifies the correct exposure region that maximizes/minimizes conditional mean and estimates the correct PIE target parameter built into a DGP for this region. Henceforth we refer to our method as CVtreeMLE.
@@ 6.1.1 Two-Dimensional Exposure Simulations
This DGP has the following characteristics, $O=(W,A,Y)$. $W$ are three baseline covariates, $W_{1}\sim\mathcal{N}(\mu=37,\sigma=3),W_{2}\sim\mathcal{N}(\mu=20,\sigma=1),W_{3}\sim\mathcal{B}(\mu=0.5)$. Here $\mathcal{B}$ is a Bernoulli distribution, and $\mathcal{N}$ is normal. These distributions and values were chosen to represent a study with covariates for age, BMI, and sex. Our generated exposures were also created to represent two chemical exposures quantized into five discrete levels. The values and range of the outcomes were chosen to represent common environmental health outcomes, such as telomere length or epigenetic expression. We are interested in sampling observations into a 2-dimensional exposure grid. We use discrete exposures to facilitate calculation of the ground truth effects without numerical integration. Here, a $5\times 5$ grid is based on combinations of two discrete exposure levels with values 1-5. We want the number of observations in each of these cells to be affected by covariates. To do this, we define a conditional categorical distribution $P\{(A_{1},A_{2})=(a_{1},a_{2})|W=w\}$ and sample from it.
$P\{(A_{1},A_{2})=(a_{k},a_{l})|W\}=\frac{e^{W^{\top}\beta_{k,l}}}{1+\sum_{k,l}e^{W^{\top}\beta_{k,l}}}$
Here, the $\beta$’s attached to each covariate were drawn from a normal distribution with means 0.3, 0.4, 0.5 and 0.5, respectively, all with a standard deviation of 2. This then gives us 25 unique exposure regions with densities dependent on the covariates. We then want to assign an outcome to each of these regions based on the main effects and interactions between exposures. We generate the outcome as: $Y=0.2A_{1}^{2}+0.5A_{1}A_{2}+0.5A_{2}^{2}+0.2*W_{1}+0.4*W_{3}+\mathcal{N}(0,0.1)$. This indicates that there is a slightly weaker squared effect for $A_{2}$ relative to $A_{1}$ and a strong interaction between the exposures and confounding due to $W_{1}$ and $W_{2}$.
@@
and select the region that matches our oracle parameter under our set of positivity restrictions. Therefore, to approximate the ARE of a region $\mathcal{A}$ to arbitrary precision we compute
$\text{ARE}(\mathcal{A})=\frac{1}{b}\sum_{i}^{b}Q_{\mathcal{A}}(1,W_{i})-Y_{i}$
using a large $b$. We repeat this for each of the $(6^{2})^{2}$ possible rectangular regions to determine which has the optimal ARE.
@@ 6.2 Evaluating Performance
For each iteration, we calculate metrics for bias, variance, MSE, CI coverage, and confusion table metrics for the true minimal region compared to the estimated region. We also calculate the bias for the estimated mean in the region found in our recursive partitioning algorithm to find the minimum region. For each type of estimate (pooled TMLE, k-fold specific TMLE estimates), we have bias when comparing our estimate to 1. The ARE based on the true region in the DGP that minimizes the outcome mean 2. The ARE when the data-adaptively determined region is applied to the population. Therefore, when comparing to the true "oracle" region ARE we have $\psi^{0}_{\text{pooled tmle MSE}}$: This is the MSE of the pooled TMLE ARE compared to the ground-truth ARE for the true oracle region built into the DGP which minimizes the mean adjusted outcome and $\psi^{0}_{\text{mean v-fold tmle MSE}}$: This is the MSE of the mean k-fold specific AREs compared to the ground-truth ARE for the true region built into the DGP.
@@ 6.4.1 CVtreeMLE Algorithm Identifies the Oracle Region with Minimum Expected Outcome
First, we describe the results to identify the true region built into the DGP. It is obviously necessary that this converges to the truth as the sample size increases in order for the estimates $\psi^{0}$ to be asymptotically unbiased. Overall, we find that our tree algorithm identifies the true region in the DGP and therefore provides results that have a minimum outcome for treatment policies. Our recursive algorithm shows very good identification of the true outcome minimizing region even in low sample sizes. At sample size 300 the average false positives were 0.053, at 500, 0.0005, and at 750 onward 0. The average true positives were 100% across all sample sizes. Our recursive partitioning algorithm correctly identifies the interaction between both exposures as the minimizing region (as opposed to marginal impacts) in 100% of folds in all sample sizes. Figure 1 shows the bias in comparing the estimated average in the estimated minimum region of the decision tree to that of the true oracle average in the oracle minimizing region. From this figure, it can be seen that we have asymptotic convergence to the true minimal average at a rate faster than $\sqrt{n}$ for this simulation.